Joe is in charge of purchasing office supplies and his budget is $175.00. Paper is $21.99 a box, toner cartridges are $52.90 each, and folders are $0.99 each. All office supplies are 20% off. If Joe needs to buy 3 boxes of paper and 2 toner cartridges, how many folders can he buy?


1
Expert's answer
2022-03-07T17:03:04-0500

3 boxes of paper and 2 toner cartridges cost "21.99(1-0.2)*3+52.90(1-0.2)*2=\$137.42."

Money left: "175.00-137.42=\$37.58."

Joe can buy: "\frac{37.58}{0.99(1-0.2)}=47" folders.